# JugglePAC: a Pipelined Accumulation Circuit

### v2

@@ V Conclusion
Accumulation is a fundamental operation, which appears in many types of computational workloads. In the case of high-throughput floating-point accumulation, complexities arise due to pipelining, especially when the system must handle consecutive datasets of varying lengths while producing results in the input order. Existing solutions often introduce significant overheads, while resulting in a reduced clock frequency or increased area complexity. In this work, we have introduced JugglePAC, a novel fully pipelined reduction circuit designed to overcome these challenges. JugglePAC leverages a single floating-point adder and yet efficiently manages accumulation tasks. Implemented and evaluated across multiple FPGAs, JugglePAC consistently delivers superior results in area and timing simultaneously, while previous works deliver either competitive area or timing, but not both at the same time.

### v3

@@ V Conclusion
Accumulation is a fundamental operation found in many computational workloads. High-throughput floating-point accumulation presents challenges due to pipelining, especially when handling consecutive datasets of varying lengths and preserving input order. Existing approaches often incur significant overhead, leading to reduced clock frequency or increased area complexity. This work introduced JugglePAC, a novel, fully pipelined reduction circuit that uses a single floating-point adder to manage accumulation efficiently. Evaluated on multiple FPGAs, JugglePAC consistently achieves both low area and high timing performance, while prior designs typically improve one but not both at the same time.